I can run 3 times faster than I can walk. One day I walked for 2 hours and ran for 30 mins.
​If I travelled 21.7 km, how fast can I walk. Use algebra to solve.

Answer by stanbon(75887)

I can run 3 times faster than I can walk. One day I walked for 2 hours and ran for 30 mins.
​If I traveled 21.7 km, how fast can I walk. Use algebra to solve.
------
Walk DATA:
distance = x km ; rate = r kph; time = x/r hrs
----
Run DATA::
distance = 21.7-x km ; rate = 3r kph ; time = (21.7-x)/(3r) hrs
-----
Equations:
x/r = 2
(21.7-x)/(3r) = 0.5
----------------------------
Substitute for "x" and solve for "r"::
(21.7-2r)/(3r) = 1/2
---------
3r = 43.4 - 4r
7r = 43.4
walk rate = 6.2 km/hr
run rate = 3(6.2) = 18.6 km/hr
---------------------
Cheers,
Stan H.
--------------